40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Entry  Thu Jul 31 15:07:53 2014, Koji, General, General, Item lending 
    Reply  Fri Jan 30 19:31:08 2015, Koji, General, General, Item lending 
       Reply  Tue Feb 3 18:23:49 2015, Koji, General, General, Item lending 
          Reply  Wed Feb 4 20:07:24 2015, Koji, General, General, Item lending C30140_1.JPGC30140_2.JPGC30141.JPGC30142.JPGC30143.JPG
             Reply  Tue Jul 21 20:20:12 2015, Koji, General, General, Item lending 
                Reply  Fri Sep 9 14:34:31 2016, Koji, General, General, Item lending 
                   Reply  Thu Feb 27 14:31:13 2020, Koji, General, General, Item lending P_20200227_134755_vHDR_On.jpg
                      Reply  Mon Aug 10 15:29:54 2020, Koji, General, General, Item lending 
Message ID: 221     Entry time: Tue Feb 3 18:23:49 2015     In reply to: 220     Reply to this: 222
Author: Koji 
Type: General 
Category: General 
Subject: Item lending 

- The laser was removed and shipped to LHO today.

- UV illuminator / fused silica fiber light guide / UV power meter / UV face shield (Qty 2) will be shipped to MIT.
They are CIT properties except for the illuminator.

Quote:

Gabriele:

PZT HV Amp

Evan:
HP signal generator (990MHz) (prev. setting 32.7MHz / +3dBm)
Black glass beam dump

Dmass:

LB1005 Oct 24.

Quote:

Tara: Laser Safety goggle -> Returned

Evan:
HP signal generator (990MHz) (prev. setting 32.7MHz / +3dBm)
Black glass beam dump

Dmass:

LB1005 Oct 24.

 

 

ELOG V3.1.3-